OpenStudy (anonymous):

hey can someone please give me COMPLETE step by step instructions on how to seperate the varialbes in this DE: dy/dx = (1+2y^2)/(ysinx)

OpenStudy (anonymous):

is this implicit differentiation

OpenStudy (anonymous):

I don't know..I'm in calc 2, were studying very basic DE's. It just says SOLVE. I know how to solve it, i just don't know how i can seperate the variables in this equation.

OpenStudy (anonymous):

Once i seperate the variables it is very easy to solve...but i dont know how to seperate the variables...

OpenStudy (anonymous):

dy/dx[1]+dy/dx[2y^2] all divided by dy/dx[ysinx]

OpenStudy (anonymous):

and your solving for y

OpenStudy (anonymous):

I dont get it...I dont need the answer, I want to learn how to seperate the variables

OpenStudy (anonymous):

thats not the answer that how you seperate them

OpenStudy (anonymous):

you will also be using the quotient rule which is fg'-gf'/g^2

OpenStudy (anonymous):

I dont understand...Can you tell me step by step what i should do, i will work out out on my paper. For example say "first you must move the y to the left by multipying both sides by y) be very specific please...i know the answer i just dont get how to get there!

OpenStudy (anonymous):

and go ONE step at a time, very slowly

OpenStudy (anonymous):

i will try and help as much as i can but am studying calc 1 and am not that good at this

OpenStudy (anonymous):

Oh ok well what i am asking is just algebraic manupulation..how can i move all the y's to one side and x's to the right

OpenStudy (anonymous):

ok so this is definitely imlicit

OpenStudy (anonymous):

yeah am stuck an sorry

OpenStudy (anonymous):

fine.. so you seperate variables get ydy/(1+2y^2) = (sinx)^-1 dx right?

OpenStudy (anonymous):

look at the left side use substitution: say u=1+2y^2 so du=4ydy but as you have y not 4y you take 1/4 outside of the integral which gives you int(du/u) which is lnu which is ln(1+2y^2)

OpenStudy (anonymous):

so the right side: integral of cscx is -ln(cot(x)+csc(x)) + C and as we have a ln on both sides we can take do e^ln(...) which negates the ln(...) on both sides

OpenStudy (anonymous):

this gives us 1+2y^2 = K(cotx +cscx) so y= sqrt(K(cotx +cscx) -1)/2)

OpenStudy (anonymous):

on second thought this may actually be totally wrong but check it for yourself